OpenStudy (anonymous):

Given the function f(x) = 0.3(4)x, what is the value of f^(−1)(6)?

OpenStudy (anonymous):

I am assuming that f^(-1) means the inverse function, and you are plugging 6 in. So to get the inverse of a function, you can switch the x and y and then solve for y. So in your example you have y=.3(4)x. Switch x and y here, and solve for y, then plug in 6 into your inverse function.

OpenStudy (anonymous):

where would 6 be plugged in

OpenStudy (anonymous):

I just did this, Step 1: replace f(x) with y. Step 2: switch x and y 3: Isolate y 4: Switch the variables back 5: Replace y with f^-1(x) 6. Plug in six for the new equation where x is. :D

OpenStudy (anonymous):

I don't know if I made it look right, but this is how it's supposed to look. f(x)=0.3(4)^x 1. y=0.3(4)^x 2.x=0.3(4)^y x/0.3=4^y log(base)4 x/0.3=log(base)4 4^y log(base)4 x/0.3=y f^-1(x)=log(base)4 x/0.3 3. Now plug in 6 for x. f^-1(x)=log(base)4 (6/0.3) f^-1(x)=log(base)4 (20) Now, solve with the change of base formula. log20/log4 Answer is roughly 2.161 (: